OpenStudy (cloverracer):

Solve the inequality. I'm stuck and want to learn how to get the answer. I've tried many ways but keep getting irrelevant answers.

OpenStudy (cloverracer):

\[57 - 4t \ge 13\]

Mehek (mehek14):

\(13-57=?\)

OpenStudy (cloverracer):

-44

Mehek (mehek14):

\(\dfrac{-44}{-4}=?\) remember - / - = +

OpenStudy (cloverracer):

How did you get -44/4 ?

Mehek (mehek14):

when you subtracted, you had \(-4t>-44\) you want to get t by itself so divide -44/-4

OpenStudy (cloverracer):

ohh okay I see this is my work..